OpenStudy (anonymous):

government of delhi allotted relief fund to help the families whose houses and shops were burnt in a fire accident.the fund is represented by 6x^3-11x^2+15x-24.the fund is equally divided between each of the families of that accident.each family receieves an amount of 3x-7.after distribution,7x+11 amount is left.the district magistrate decided to use this amount to develop the infrastructure of the area.find the number of families which received relief fund from government.what value is depucted here

OpenStudy (anonymous):

@amistre64 @satellite73

OpenStudy (amistre64):

so, 11x^2 +15x -24 is equal to (3x-7) Q(x) + 7x+11

OpenStudy (amistre64):

6x^3 went missing lol

OpenStudy (amistre64):

for simplicity: P = LQ + R P/L = Q + R/L P/L - R/L = Q

OpenStudy (amistre64):

\[Q(x)=\frac{(6x^3-11x^2+15x-24)-(7x+11)}{3x-7}\]

OpenStudy (anonymous):

can u suggest a value???

Parth (parthkohli):

Boo, value-based questions. The values depicted by the government of Delhi are: - Helpfulness. - Obligation. - Responsibility.

OpenStudy (anonymous):

Thanks buddy....